# Rethink the Evaluation Protocol of Model Merging on Classification Task

### v1

@@ Abstract
Model merging combines multiple fine-tuned models into a single one via parameter fusion, achieving improvements across many tasks. However, in the classification task, we find a misalignment issue between merging outputs and the fine-tuned classifier, which limits its effectiveness. In this paper, we demonstrate the following observations: (1) The embedding quality of the merging outputs is already very high, and the primary reason for the differences in classification performance lies in the misalignment issue. (2) We propose FT-Classifier, a new protocol that fine-tunes an aligned classifier with few-shot samples to alleviate misalignment, enabling better evaluation of merging outputs and improved classification performance. (3) The misalignment is relatively straightforward and can be formulated as an orthogonal transformation. Experiments demonstrate the existence of misalignment and the effectiveness of our FT-Classifier evaluation protocol.
Rethink the Evaluation Protocol of Model Merging on Classification Task
@@ 1 Introduction
Fine-tuning a pre-trained model (PLM) on downstream tasks to obtain a task-specific fine-tuned model is a common practice nowadays Devlin et al. (2018). Building on this training paradigm, real-world applications increasingly aim to leverage a single model that can be fine-tuned across multiple tasks, thus enabling it to address all tasks simultaneously Zhang and Yang (2021). This is known as multi-task learning (MTL), which generally aggregates all task data for joint training. However, MTL not only requires the full amount of task data, which raises data privacy concerns, but also involves a fully fine-tuning process, leading to high training costs and poor scalability Yang et al. (2024b). Therefore, researchers are actively exploring model merging as an alternative Izmailov et al. (2018); Jin et al. (2022); Ilharco et al. (2022); Yang et al. (2024c, a).
@@
We argue that this is because classification tasks typically require a classifier (e.g., the final layer of a neural network) for evaluation. However, since the classifier’s weight dimensions vary across tasks (due to different numbers of classes), classifiers cannot be merged directly. As a result, evaluation of the merged model depends on the fine-tuned classifier, causing a misalignment between the merging output and the classifier, which ultimately leads to suboptimal performance in classification tasks.
To support our argument, we visualize the merging outputs of three representative methods on text classification datasets using T-SNE Van der Maaten and Hinton (2008), as shown in Figure 1. From the figure, we observe two key phenomena:
Refer to caption Refer to caption
(a) Weight Averaging (79.2%) (b) Weight Averaging (55.1%)
Refer to caption Refer to caption
(c) Task Arithmetic (82.9%) (d) Task Arithmetic (81.5%)
Refer to caption Refer to caption
(e) Ties (83.3%) (f) Ties (83.4%)
Figure 1: T-SNE visualization of three representative merging methods on AG News and DBpedia. The blue points represent the sentence embeddings of the fine-tuned model, while the red points represent the results of the merging model. The color shades indicate different labels. The value indicates the classification accuracy of merging methods. For results on other datasets, please refer to Appendix A.1.
(1) Representation bias between merging and fine-tuned models. The overall red and blue clusters are non-overlapping, indicating that merging outputs and fine-tuned model outputs exhibit a bias in the latent space. This phenomenon is consistent with the observations in Yang et al. (2024b).
(2) Good clustering effects in merging outputs. Despite notable gaps in classification accuracy between different merging methods (e.g., nearly a $30\%$ accuracy difference between Weight Averaging and Ties on DBpedia), their merging outputs exhibit similar clustering performance. This suggests that merging outputs retain high-quality embeddings, comparable to fine-tuned models.
These observations lead us to suspect reasonably: There is a representation misalignment issue between merging outputs and the fine-tuned classifier, and the current evaluation protocol appears incapable of accurately assessing the true quality of merging methods. Therefore, in this work, we conduct the following analysis to validate our suspicion.
Assessing the quality of merging outputs
@@
We bypass the classifier and instead use a few-shot set of labeled samples as class anchors. Employing a KNN-based method—commonly applied in sentence representation studies—we directly measure classification precision. The results are reported in Figure 2. Surprisingly, the KNN-based evaluation outperforms the current evaluation protocol on three representative methods, even with limited labeled data, demonstrating the high quality of the merging outputs. Moreover, this also indicates that the performance drop of the current evaluation protocol mainly contributes to the misalignment issue.
Alignment formulation
We designed two fine-tuning strategies: (1) adding a mapping matrix between the merging output and the fine-tuned classifier, and (2) directly fine-tuning the classifier to adapt to the merging output. The results are shown in Figure 4. The findings confirm that aligning either the merging output or the classifier can significantly improve classification performance, with results even approaching the upper limit (the fine-tuned model’s performance).
New evaluation protocol for misalignment alleviation
@@
Based on the above observations, we believe that the current evaluation protocol mainly assesses the compatibility between the merging output and the fine-tuned classifier rather than effectively evaluating the quality of the merging outputs themselves, which in turn limits the final classification performance. Therefore, we propose the FT-Classifier evaluation protocol, which leverages few-shot unlabeled data and employs minimal training steps without introducing additional parameters, aiming to train an aligned classifier and then make the evaluation more convincing. The results of the FT-Classifier evaluation protocol, presented in Figure 5, demonstrate that only a minimal amount of data is required for the alignment, resulting in a significant improvement in classification performance.
The misalignment is an orthogonal transformation
The fact that the misalignment issue can be mitigated with only a small amount of data demonstrates that the transformation of misalignment is exceedingly straightforward. Intuitively, we hypothesize that this transformation can be approximated as a sufficiently simple orthogonal transformation. To support this, we add an orthogonal transformation regularization to the mapping matrix. The results shown in Table 2 represent that it is possible to train an orthogonal transformation with a slight drop in classification performance. This means that, through a rigid transformation such as rotation or reflection, the merging outputs can be aligned with the fine-tuned classifier. It indicates that the merging outputs, compared to the fine-tuned model’s outputs, preserve the intrinsic properties of the manifold, which are essential for classification.
In summary, we make the following contributions:
@@
  • Through empirical analysis, we reveal that the misalignment issue between merging outputs and the fine-tuned classifier renders the current evaluation protocol ineffective for accurately assessing merging methods.
  • We introduce a new FT-Classifier evaluation protocol that alleviates the misalignment issue by limited data, enabling a more accurate assessment of merging methods while significantly enhancing their performance.
@@
  • We prove that this misalignment can converge to an orthogonal transformation, indicating that the intrinsic properties of the manifolds in the merging outputs remain consistent with those of the fine-tuned model.
@@ 2.1 Definition of Model Merging
Let $f_{b}$ denote the pre-trained base model with parameter $\mathbf{\Theta}_{b}$. Given $T$ specific tasks with training data $X_{t}$ and corresponding label $Y_{t}$ for each, one can obtain $T$ fine-tuned models. And the fine-tuned model for task $t$ can be denoted as $f_{t}$ with parameter $\mathbf{\Theta}_{t}$. Without fully fine-tuning, model merging aims to merge these homologous models to generate a single $f_{m}$ with parameter $\mathbf{\Theta}_{m}$, which can simultaneously perform well across all $T$ tasks.
@@ 2.2 Representative Methods
Up to date, numerous model merging approaches have been proposed, which we will discuss in detail in Section 7. In this section, we primarily focus on introducing the three most representative methods. Weight averaging Wortsman et al. (2022) simply averages $T$ fine-tuned models for merging, which is formulated as:
@@
2.3 Current Evaluation Protocol
@@
Existing merging methods primarily focus on generating a comprehensive encoder that absorbs the ability from each fine-tuned model. While the merging encoders $f_{m}$ are evaluated on the fine-tuned classier, which can be formulated as:
$\hat{Y}_{t}=f_{m}(X_{t};\mathbf{\Theta}_{m})\cdot\mathbf{W}_{t},$ (4)
@@
where $\mathbf{W}_{t}$ represents the weight of the classifier of task $t$. It maps the merging outputs to the predicted labels $\hat{Y}_{t}$.
3 Existence of the Misalignment Issue
@@
From Eq. 4, we can easily find the fact that due to the change in parameters, the merging outputs of $X_{t}$ generated by $f_{m}$ are inevitably shifted in the latent space compared to those generated by $f_{t}$. However, the evaluation of $f_{m}$ still relies on $\mathbf{W}_{t}$, which is tailored to $f_{t}$. Naturally, this raises a question: Does the outputs of $f_{m}$ align well with the fine-tuned classifier $\mathbf{W}_{t}$?
3.1 Merging outputs Quality Evaluation
As shown and analyzed in Figure 1, we speculate that the merging outputs produced by different $f_{m}$ consistently exhibit high quality. To validate our hypothesis, we use a KNN-based method to bypass the classifier and directly evaluate the classification performance of different merging outputs. Illustrations of the current evaluation and this KNN-based one are presented in Figure 3 (a) and (b).
@@ 3.1 Merging outputs Quality Evaluation
Specially, for the KNN evaluation, we first randomly select few-shot labeled data as anchors for each class, then calculate the average Euclidean Distance between the test sample and these anchors for each class, and finally assign the class with the minimum distance as the prediction. The overall process is formulated as:
$\hat{y}_{t}=\arg\min_{y_{t}\in Y_{t}}\frac{1}{k}\sum_{i=1}^{k}d\left(X_{t},[A_{t}^{y_{t}}]^{i}\right).$ (5)
@@
where $y_{t}$ is an element of $Y_{t}$, $k$ refers to the number of few-shot samples, $d$ represents the Euclidean Distance metric, $A_{t}^{y_{t}}$ represents the randomly selected anchors for label $y_{t}$, and []delimited-[][$\cdot$][ ⋅ ] denotes the indexing function. From Eq. 5, we observe that the KNN evaluation directly utilizes merging outputs for classification with only few-shot labeled data, and we believe this approach provides an effective assessment of merging outputs quality.
Refer to caption Refer to caption Refer to caption
@@
Figure 2: The performance of KNN evaluation compared with current evaluation protocol under varying numbers of few-shot examples $k$. The $x$-axis represents the $k$, and the $y$-axis represents the average of Accuracy for all merging tasks. Experimental details and explanations are provided in Section 6.
@@
Results of KNN evaluation compared with the current evaluation protocol and upper limits fine-tuned models are shown in Figure 2. From the figure, we observe that directly using the KNN evaluation yields excellent classification results. Even with only 5 few-shot examples, the KNN evaluation outperforms the current protocol, which relies on a fine-tuned classifier. This suggests that the outputs of different merging methods are of high quality, and the poor performance in the current evaluation protocol is more likely caused by the misalignment between the merging outputs and the fine-tuned classifier.
3.2 Revisiting the Objective of Model Merging
Revisiting the target of model merging, which aims to generate a new model $f_{m}$ using $f_{1},...,f_{T}$ that performs well across all $T$ tasks, we note that the upper bound of $f_{m}$’s performance on a specific task $t$ corresponds to the result from fine-tuned model $f_{t}$. Naturally, the objective of model merging is to minimize this performance discrepancy of $f_{m}$ and $f_{t}$ Ortiz-Jimenez et al. (2024), which is defined as:
$\arg\min_{\mathbf{\Theta}_{m}}\sum_{t=1}^{T}g(f_{m}(X_{t};\mathbf{\Theta}_{m})\cdot\mathbf{W}_{t},f_{t}(X_{t};\mathbf{\Theta}_{t})\cdot\mathbf{W}_{t})$ (6)
where $\cdot$ represents matrix multiplication, and $g$ is a measurement function.
From the objective, we observe that the optimal $\mathbf{\Theta}_{m}^{*}$ should not only be related to the fine-tuned model $\mathbf{\Theta}_{t}$ but also to $\mathbf{W}_{t}$. Since we have also observed that the merging output and classifier exhibit a certain misalignment, we hypothesize that alleviating the misalignment in latent space could lead to a better evaluation of $f_{m}$.
@@ 3.2 Revisiting the Objective of Model Merging
3.3 Alignment Formulation
Consequently, we conduct two simple strategies to align merging outputs and the fine-tuned classifier: (1) introducing a mapping function between the outputs and the classifier to learn the alignment, and (2) directly fine-tuning the classifier to better align with the outputs. Both training approaches used Eq. 6 as the objective, with KL divergence Kullback and Leibler (1951) as the metric function. Following Yang et al. (2024b), we employ unlabeled test data to train the related parameters.
Specifically, for the first approach, we simply define the mapping function as a square matrix $\mathbf{M}_{t}$ with the same dimensionality as the hidden layer, and the optimization objective, named Aligned $\mathbf{M}_{t}$ is defined as follows:
[…]
Note that during the training process, only $\mathbf{M}_{t}$ is updated, while all other parameters remain frozen.
For the second approach, we use $\mathbf{W}_{t}$ to initialize a new classifier $\mathbf{W}_{t}^{\prime}$, and define the optimization objective Aligned $\mathbf{W}_{t}^{\prime}$ as follows:
$\arg\min_{\mathbf{W}_{t}^{\prime}}\sum_{t=1}^{T}{\rm KL}(f_{m}(X_{t};\mathbf{\Theta}_{m})\!\cdot\!\mathbf{W}_{t}^{\prime},f_{t}(X_{t};\mathbf{\Theta}_{t})\!\cdot\!\mathbf{W}_{t})$ (8)
In this objective, only the new classifier $\mathbf{W}_{t}^{\prime}$ requires updating.
Figure 4: Comparison of the Current Eval and two Aligned variants.
Illustrations of Aligned $\mathbf{M}_{t}$ and Aligned $\mathbf{W}_{t}^{\prime}$ are shown in Figure 3 (c), (d), respectively. The relevant comparable results are shown in Figure 4. From the Figure, we observe that:
(1) With only a limited number of parameters, the misalignment can be captured, achieving surprising classification performance approaching the upper bound of the fine-tuned model. Both aligned variants require very few parameters, with the Aligned $\mathbf{W}_{t}^{\prime}$ variant even using the exact same number of parameters as the fine-tuned classifier $\mathbf{W}_{t}$. This also indicates that the quality of outputs via different merging methods is already high.
(2) The quality of the outputs produced by different $f_{m}$ does not differ significantly, and the current evaluation protocol does not assess $f_{m}$ well. Since the performance gap of different aligned variants is much smaller than the current evaluation of $f_{m}$. For example, the performance of Weight Averaging was initially around 10% lower than that of Task Arithmetic, but after alignment, their performances became nearly identical.
These findings demonstrate that the misalignment issue hinders the current evaluation protocol from effectively assessing the capacity of $f_{m}$. And alleviating the misalignment problem could potentially improve classification performance.
Method AG News Yelp Amazon Yahoo DBpedia Avg.
FT 91.6 60.4 57.9 72.7 98.8 76.3
MTL 90.6 59.1 55.6 71.3 98.5 75.0
WA 79.2 49.8 45.0 50.3 55.1 55.8
+FT-C 85.9 49.8 46.8 66.1 96.5 69.0
TA 82.9 55.8 48.4 53.1 81.5 64.3
+FT-C 87.5 57.0 52.5 67.9 96.6 72.3

### v2

# Fine-tuning Aligned Classifiers for Merging Outputs: Towards a Superior Evaluation Protocol in Model Merging

@@ Abstract
Model merging combines multiple fine-tuned models into a single one via parameter fusion, achieving improvements across many tasks. However, in the classification task, we find a misalignment issue between merging outputs and the fine-tuned classifier, which limits its effectiveness. In this paper, we first demonstrate the following observations: (1) Merging outputs exhibit the comparable cluster effect with fine-tuned outputs, and already contain necessary classification information; (2) The misalignment between merging outputs and the fine-tuned classifier can converge to an orthogonal transformation, and alleviating this misalignment can significantly enhance the performance of merging models. Based on these observations, we then propose a new protocol FT-Classifier, which fine-tunes an aligned classifier with few-shot unlabeled samples, enabling better evaluation of merging methods and improved classification performance.
Fine-tuning Aligned Classifiers for Merging Outputs:
Towards a Superior Evaluation Protocol in Model Merging
@@ 1 Introduction
Fine-tuning a pre-trained language model on downstream tasks to obtain a task-specific fine-tuned model is a common practice nowadays Devlin et al. (2018). Building on this training paradigm, real-world applications increasingly aim to leverage a single model that can be fine-tuned across multiple tasks, thus enabling it to address all tasks simultaneously Zhang and Yang (2021). This is known as multi-task learning (MTL), which generally aggregates all task data for joint training. However, MTL not only requires the full amount of task data, which raises data privacy concerns, but also involves a fully fine-tuning process, leading to high training costs and poor scalability Yang et al. (2024b). Therefore, researchers are actively exploring model merging as an alternative Izmailov et al. (2018); Jin et al. (2022); Ilharco et al. (2022); Yang et al. (2024c, a).
@@
We argue that this is due to the inevitable changes in the parameters of the merged model compared to the fine-tuned model. However, the current evaluation protocol (Current Eval) still utilizes the fine-tuned classifier for merged model assessment. This misalignment between merging outputs and the fine-tuned classifier leads to inaccurate evaluations of the merging method. In this work, we conduct the following analysis to validate our argument.
Merging outputs already exhibit a good clustering effect.
We visualize the merging outputs of three representative methods on text classification datasets using T-SNE Van der Maaten and Hinton (2008), as shown in Figure 1. From the Figure, we observe that although the performance of different merging methods varies across specific tasks, all merging outputs of representative methods exhibit comparable clustering performance to that of fine-tuned outputs. This motivates us to ponder that the merging outputs already include the necessary classification information, yet it is not accurately evaluated in the Current Eval.
Misalignment between merging outputs and fine-tuned classifier leads to the undervaluation of merging methods.
@@
We bypass the fine-tuned classifier and instead use a few-shot set of labeled samples as class anchors. Employing a KNN-based method—commonly applied in sentence representation studies—we directly measure classification accuracy. The results of this KNN evaluation protocol (KNN Eval) are reported in Figure 2. Surprisingly, KNN Eval significantly outperforms the Current Eval on three representative methods, even with limited labeled data, demonstrating that Current Eval tends to underestimate merging methods. Since KNN Eval does not require a classifier at all, while Current Eval requires a fine-tuned classifier, this also proves that this underestimation may be caused by the misalignment between merging outputs and the fine-tuned classifier.
The misalignment can converge into an orthogonal transformation.
In conjunction with Figure 1, which illustrates a certain bias between merging outputs and fine-tuned outputs in the latent space, we hypothesize some simulation of transformation to align them. Specifically, we employ a series of transformations, varying the strength of the constraints, for empirical verification. The definition of different transformations and corresponding results are detailed in Table 1 and Figure 4, respectively. The results demonstrate that after transformation alignment, all merging methods exhibit significant improvement compared to the Current Eval, even approaching the upper bound of the fine-tuned model. Moreover, when these transformations converge to an orthogonal transformation, there is only a slight drop in classification performance. This indicates that through a rigid transformation such as rotation or reflection, the misalignment can be easily alleviated. As a result, the fine-tuned classifier may only need a simple alignment with the merging outputs to achieve better evaluation results.
Fine-tuning an aligned classifier can alleviate the misalignment and facilitate merging methods assessment.
@@
Based on the above observations, we believe that Current Eval mainly assesses the compatibility between the merging outputs and the fine-tuned classifier rather than effectively evaluating the capability of the merging outputs themselves, which in turn limits the final classification performance. Therefore, we propose the FT-Classifier evaluation protocol (FT-Classifier Eval), which leverages few-shot unlabeled data to train an aligned classifier for merging outputs. This protocol does not introduce additional parameters and requires only minimal training steps, making it a convincing alternative to Current Eval. The results of the FT-Classifier Eval, as presented in Figure 5, demonstrate that such a lightweight and computationally feasible post-processing operation improves the effectiveness of merging methods assessment.
In short, we make the following contributions:
@@
  • Through empirical analysis, we reveal that the misalignment issue between merging outputs and the fine-tuned classifier renders the Current Eval ineffective for accurately assessing merging methods.
@@
  • We prove that this misalignment can converge to an orthogonal transformation, indicating that the intrinsic structure and geometric properties of the merging outputs remain consistent with those of the fine-tuned model.
  • We introduce a new FT-Classifier Eval that alleviates the misalignment issue using limited data, without adding parameters or changing the model structure. This enables a more accurate assessment of merging methods while significantly enhancing their performance.
@@ 2.1 Definition of Model Merging
Let $f_{b}$ denote the pre-trained base model with parameter $\mathbf{\Theta}_{b}$. Given $T$ specific tasks with training data $X_{t}$ and corresponding label $Y_{t}$ for each data, one can obtain $T$ fine-tuned models. The fine-tuned model for task $t$ can be denoted as $f_{t}$ with parameter $\mathbf{\Theta}_{t}$. Without fully fine-tuning, model merging aims to merge these homologous models to generate a single $f_{m}$ with parameter $\mathbf{\Theta}_{m}$, which can simultaneously perform well across all $T$ tasks.
@@ 2.2 Representative Methods
Up to date, numerous model merging approaches have been proposed, which we will discuss in detail in Section 7. In this section, we primarily focus on introducing the three most representative methods.
Weight averaging Wortsman et al. (2022) simply averages $T$ fine-tuned models for merging, which is formulated as:
@@
2.3 Evaluation of Merging Models
@@
Existing merging methods in classification tasks primarily focus on merging fine-tuned encoders, while not performing the merging operation on fine-tuned classifiers. That is, the merging encoders $f_{m}$ should be evaluated on the fine-tuned classier, which can be formulated as:
$\hat{Y}_{t}=f_{m}(X_{t};\mathbf{\Theta}_{m})^{\top}\mathbf{W}_{t},$ (4)
@@
where $\mathbf{W}_{t}$ represents the weight of the fine-tuned classifier of task $t$, which maps the merging outputs to the predicted labels $\hat{Y}_{t}$. In the following descriptions, this currently adopted evaluation protocol is referred to as ‘Current Eval’.
3 Existence of the Misalignment
@@
From Eq. 4, we can easily find the fact that due to the change in parameters, the merging outputs of $X_{t}$ generated by $f_{m}$ are inevitably shifted in the latent space compared to those generated by $f_{t}$. However, the evaluation of $f_{m}$ still relies on $\mathbf{W}_{t}$, which is tailored to $f_{t}$. Naturally, this raises a question: Does the outputs of $f_{m}$ align well with the fine-tuned classifier $\mathbf{W}_{t}$? To analyze this issue, this section delves into an in-depth analysis of the output of $f_{m}$.
3.1 Visualization of Merging outputs
Refer to caption Refer to caption
(a) Weight Averaging (79.2%) (b) Weight Averaging (55.1%)
Refer to caption Refer to caption
(c) Task Arithmetic (82.9%) (d) Task Arithmetic (81.5%)
Refer to caption Refer to caption
(e) Ties (83.3%) (f) Ties (83.4%)
Figure 1: T-SNE visualization of three representative merging methods on AG News and DBpedia. The blue points represent the sentence embeddings of the $f_{t}$ model, while the red points represent the results of the $f_{m}$ model. The color shades indicate different labels. The value indicates the classification accuracy of merging methods. For results on other datasets, please refer to Appendix A.1.
Motivated by Yang et al. (2024b), we used T-SNE to visualize the outputs of $f_{t}$ and three representative $f_{m}$ models. Furthermore, to observe the spatial variations of $f_{m}$’s output in more detail, we incorporated the corresponding label information into the visualization process, as shown in Figure 1. From this Figure, we observe two key phenomena:
(1) There is a representation bias between the merging outputs and the fine-tuned outputs. The overall red and blue clusters are non-overlapping, indicating that merging outputs and fine-tuned outputs exhibit a bias in the latent space. This phenomenon is consistent with the observations in Yang et al. (2024b).
(2) All merging outputs exhibit good clustering effects like fine-tuned outputs. Despite notable gaps in classification accuracy between different merging methods (e.g., nearly a $30\%$ accuracy difference between Weight Averaging and Ties on DBpedia), their merging outputs exhibit similar clustering performance. This suggests that all merging outputs retain necessary classification information, comparable to fine-tuned outputs.
Based on these two phenomena, we hypothesize that the misalignment between merging outputs and the fine-tuned classifier leads to an underestimation of merging methods in Current Eval.
3.2 KNN Evaluation of Merging outputs
@@
To validate our hypothesis, we use a KNN-based method to directly evaluate the classification performance of different merging outputs. Specially, for the KNN evaluation (KNN Eval), we first randomly select few-shot labeled data as anchors for each class, then calculate the average Euclidean Distance between the test sample and these anchors for each class, and finally assign the class with the minimum distance as the prediction. The overall process is formulated as:
$\hat{y}_{t}=\arg\min_{y_{t}\in Y_{t}}\frac{1}{k}\sum_{i=1}^{k}d($ $f_{m}(X_{t};\mathbf{\Theta}_{m})),$ (5)
$f_{m}([A_{t}^{y_{t}}]^{i};\mathbf{\Theta}_{m}))).$
@@
where $y_{t}$ is an element of $Y_{t}$, $k$ refers to the number of few-shot samples, $d$ represents the Euclidean Distance metric, $A_{t}^{y_{t}}$ represents the randomly selected anchors for label $y_{t}$, and []delimited-[][$\cdot$][ ⋅ ] denotes the indexing function. Illustration of Current Eval and KNN Eval are shown in Figure 3 (a) and (b).
@@
Figure 2: The performance of KNN Eval compared with Current Eval under varying numbers of few-shot examples $k$. Fine-tuned represents the results obtained using $f_{t}$, which serves as the evaluation upper bound. The $x$-axis represents the $k$, and the $y$-axis represents the average Accuracy for all merging tasks. Corresponding detailed results while $k=5$ are shown in Table 2. Experimental details and explanations are provided in Section 6.
From Eq. 5, we observe that KNN Eval directly utilizes the merging outputs for classification with only a few labeled samples. We believe this evaluation protocol can effectively assess whether the necessary classification information is embedded in the merging outputs. Results of KNN Eval compared with Current Eval are shown in Figure 2.
@@
From the Figure, we observe that KNN Eval surprisingly yields excellent classification results. Even with only 5 few-shot examples, the KNN Eval outperforms the Current Eval, which relies on a fine-tuned classifier. This demonstrates that the outputs of different merging methods contain sufficient information for classification. Therefore, the reason for the poor performance in the Current Eval is a misconception, more likely caused by the misalignment between the merging outputs and the fine-tuned classifier.
@@
4 Analysis of the Misalignment
Through empirical analysis, we observe that there exists a misalignment issue between the merging outputs and the fine-tuned classifier, and this misalignment leads to the undervaluation of $f_{m}$’s capability in the Current Eval. Therefore, we hypothesize that: Alleviating the misalignment in latent space could lead to a better evaluation of $f_{m}$.
4.1 Different Transformation Hypothesis
To further explore this misalignment, we hypothesize a transformation between the merged outputs and the fine-tuned outputs, and we propose conjectures based on varying degrees of constraint on this transformation. Specifically, we align the merged outputs with the fine-tuned outputs by optimizing the following objective function:
$\arg\min_{G_{t}}||G_{t}(f_{m}(X_{t};\mathbf{\Theta}_{m}))-f_{t}(X_{t};\mathbf{\Theta}_{t})||_{1},$ (6)
where $G_{t}(\cdot)$ represents the transformation for task $t$, which can be instantiated using the transformations listed in Tabel 1. In particular, we employ unlabeled test data to train the related parameters. Noted, during the training process, only parameters in $G_{t}(\cdot)$ are updated, while all other parameters remain frozen. We refer to these merging outputs aligned method as ‘Aligned $f_{m}$’, illustrated in Figure 3 (c).
$G(\mathbf{x})=\mathbf{A}\mathbf{x}+\mathbf{b}$, where $\mathbf{A}\in\mathbb{R}^{n\times n},\mathbf{b}\in\mathbb{R}^{n}$
Transformation Constraints
Affine Transform /
Linear Transform $\mathbf{b}=0$
Invertible Linear Transform $\det(\mathbf{A})\neq 0,\mathbf{b}=0$
Euclidean Transform $\mathbf{A}^{\top}\mathbf{A}=\mathbf{I},\mathbf{b}\neq 0$
Orthogonal Transform $\mathbf{A}^{\top}\mathbf{A}=\mathbf{I},\mathbf{b}=0$
Scaling Transform $\mathbf{A}={\rm diag}(\mathbf{A}),\mathbf{b}=0$
Table 1: Definition of different transformations. When $\mathbf{x}$ is instantiated as the merging output, $G(\mathbf{x})$ represents the fine-tuned output. From Affine Transform to Scaling Transform, the constraints gradually become stronger.
Taking the orthogonal transformation as an example, we first randomly initialize a mapping matrix $\mathbf{A}_{t}$ and set $\mathbf{b}_{t}=0$ for each task $t$. Then, we perform optimization by imposing regularization constraints on the mapping matrix $\mathbf{A}_{t}$. The overall loss function is defined as follows:
$\arg\min_{\mathbf{A}_{t}}$ $||\mathbf{A}_{t}f_{m}(X_{t};\mathbf{\Theta}_{m})-f_{t}(X_{t};\mathbf{\Theta}_{t})||_{1}$ (7)
$+\alpha||\mathbf{A}_{t}^{\top}\mathbf{A}_{t}-\mathbf{I}||_{1},$
where $\alpha$ is the weight hyperparameter, and $\mathbf{I}$ is an identity matrix. By adjusting the value of $\alpha$, $\mathbf{A}_{t}$ can converge to an orthogonal transformation.
Results of all these Aligned $f_{m}$ are shown in Figure 4. From the Figure, we observe a very interesting phenomenon: When the transformation constraints are relaxed, the classification performance of all $f_{m}$ is barely affected and remains close to the upper bound (Fine-tuned results). For example, when the constraints are weaker than the orthogonal transformation, the performance is essentially unaffected, with at most nearly a 2% drop. This demonstrates that through an orthogonal transformation, the merging outputs can be aligned well with the fine-tuned outputs.
Figure 4: Comparison of different transformations. WA represents Weight Averaging, TA represents Task Arithmetic.
Method AG News Yelp Amazon Yahoo DBpedia Average
Fine-tuned model 91.6 60.4 57.9 72.7 98.8 76.3
Multi-task Learning 90.6 59.1 55.6 71.3 98.5 75.0
Weight Averaging 79.2 49.8 45.0 50.3 55.1 55.8
+ KNN Eval 81.9 $\pm$ 1.6 42.1 $\pm$ 4.3 41.2 $\pm$ 1.5 54.1 $\pm$ 3.5 91.5 $\pm$ 1.8 62.2 $\pm$ 0.6
+ FT-Classifier Eval 86.0 $\pm$ 1.0 51.5 $\pm$ 1.9 46.3 $\pm$ 0.6 66.2 $\pm$ 1.1 95.8 $\pm$ 0.4 69.2 $\pm$ 0.4
Task Arithmetic 82.9 55.8 48.4 53.1 81.5 64.3
+ KNN Eval 86.4 $\pm$ 1.3 54.2 $\pm$ 2.6 49.0 $\pm$ 1.6 61.6 $\pm$ 1.0 96.3 $\pm$ 0.2 69.5 $\pm$ 1.2
+ FT-Classifier Eval 87.5$\pm$ 0.7 55.6$\pm$ 0.5 51.5 $\pm$ 0.6 67.6 $\pm$ 0.7 96.6 $\pm$ 0.3 71.8 $\pm$ 0.2